OpenStudy (anonymous):

Find this number to the nearest tenth. What number is 11 1/3 of 215?

OpenStudy (mathstudent55):

Can you write this quastion in words? What number is eleven and a third of 215? Is it eleven and one third percent? I don't know what you mean.

OpenStudy (anonymous):

Yes,It's 11 1/3 %

OpenStudy (mathstudent55):

You can use a proportion: 215 is 100% x is 11 1/3% 215/100 = x/(11 1/3) Cross multiply: 100x = 215(11 1/3), but 11 1/3 = 34/3 so 100x = 215 x 34/3 100x = 7310/3 x = 7310/300 Simplify x = 731/30

OpenStudy (anonymous):

So it's 24.36 So should I round it to 24? or 24.4?

OpenStudy (anonymous):

I think it's 24.4 since it's the nearest tenth. Right?

OpenStudy (mathstudent55):

yes

OpenStudy (anonymous):

Okay.Thank you.:)